Ticket Number: HEA-NOC/20080515-2 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080515 08:53 UTC+1 Problem Start: 20080515 07:30 UTC+1
Ticket Update: 20080515 09:23 UTC+1 Problem End: 20080515 09:00 UTC+1
Migration of BCFE's current connections to HEAnet to new IP Network
Service interruption is expected due to repatching cables from the access
router in Citywest to CRS router, BGP sessions need to be reset and
will be done gradually. A side effect of this migration is asymmetrical
routing while convergence takes place.
Once the changes are completed, traffic will use the new connections,
continuing to have the current ones as backup. Eventually these will be
Contingency plan is to revert changes to existing setup by disabling any
routing over the new links.
20080514 11:00 GL Confirmed with BCFE that migration would take place
20080515 at 07:30.
20080515 08:30 GL Migration completed. BCFE's traffic is now routing via
Citywest with Kilcarbery as backup. Will monitor for next few days
20080515 09:21 BN Duplex mismatches are being recorded on this link.
Smart have been contacted and BCFE informed that troubleshooting may cause
a brief outage today to the circuit.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080515-2
HEAnet Limited. Registered in Ireland, No. 275301.